OpenStudy (anonymous):

~^.^~ check my work pweeease medal + fan questions below they might be wrong because tbh I'm not good in maths so please don't be harsh :/

OpenStudy (johnweldon1993):

Cant be harsh? -Sigh- No jk XD Okay lets start with 1) I get a different answer Completing the square tells us to take the coefficient of 'x' (10) divide it by 2 and then square that... so 10/2 = 5 and 5^2 = 25 <--

OpenStudy (anonymous):

lol xP and OMFGGGGG thanks for the explanation.

OpenStudy (johnweldon1993):

Lol of course :D Alright so hmm, let me check #2 *I'll take a bit longer to make you worry ;P

OpenStudy (anonymous):

Hey! lol xP Oh man I wish my math teacher had your sense of humor it would be a lot more fun to do.

OpenStudy (johnweldon1993):

So #2 I get something different again...so here I'm actually gonna go through the whole thing and make sure you got the process down :)

OpenStudy (johnweldon1993):

So we have \(\large x^2 - 8x + 5 = 0\) So to complete the square...we need this in the form \(\large a^2 - bx = c\) So we need to subtract 5 from both sides of this equation to make ours look like that \[\large x^2 - 8x = -5\] Good so far? :)

OpenStudy (anonymous):

Yup got it so far :]

OpenStudy (johnweldon1993):

Cool so now...from that general equation \[\large ax^2 - bx = c\] **notice i forgot the 'x' up there in the previous post...oops sorry :P Okay but from here, to complete the square, we take the coefficient of the 'x' *which is 'b' here* ...divide it by 2...and then square the result So in your equation \[\large x^2 - \color \red{8}x = -5\] 8 is the coefficient we want to work with...so 8 divided by 2 = 4....and 4 squared = 16 And with that 16...we add that to both sides of our equation \[\large x^2 - 8x + 16 = -5 + 16\] Still good? *stop me if you need it, I'm here for ya xP

OpenStudy (anonymous):

Ohhhhhhhhhh oh my I think I'm starting to see what I did wrong. Still good! :]

OpenStudy (johnweldon1993):

Awesome :D Okay so now lets simplify a bit \[\large x^2 - 8x + 16 = 11\] right? We can write that left hand side a little simpler right? \[\large (x - 4)^2 = 11\] Are you good with that?

OpenStudy (anonymous):

So it's either the first or the second answer? and yup I'm good with that. :]

OpenStudy (johnweldon1993):

Haha oh jumping the gun huh ;P Lets finish up \[\large (x - 4)^2 = 11\] Now we just take the square root of both sides \[\large \sqrt{(x - 4)^2} = \sqrt{11}\] Becauseeeeeee that cancels out the \[\large (x - 4)\cancel{^2}\] So now we have \[\large x - 4 = \pm\sqrt{11}\] And finally that gives ussssssssss....?

OpenStudy (anonymous):

The first one? O.O

OpenStudy (johnweldon1993):

Hey hey hey no! :P Solve for 'x' \[\large x - 4 = \sqrt{11}\] Add 4 to both sides \[\large x = 4 \pm \sqrt{11}\]

OpenStudy (anonymous):

Ohhhh omg darn but thanks for explaining! xP

OpenStudy (johnweldon1993):

Haha well I mean...you were close ;D had a 50 50 chance :P

OpenStudy (anonymous):

Haha thanks!

OpenStudy (johnweldon1993):

For 3) ...I seem to get a different answer again Sheesh I need a break, lets get dinner quick ;D XDDDDD jk jk haha okay 3 \[\large f(x) = 2(x - 1)^2 + 3\] WE need to focus on exponents first here \[\large 2(x - 1)(x - 1) + 3\] Lets do all the multiplication next \[\large 2(x^2 - 2x + 1) + 3\] \[\large 2x^2 - 4x + 2 + 3\] \[\large 2x^2 - 4x + 5\]

OpenStudy (anonymous):

Oh man I would love some dinner count me in ^.^ Ah oh gosh I'm seeing so many of my mistakes lol but I really appreciate you correcting and showing me how it's really done.

OpenStudy (johnweldon1993):

Awesome, reservations are at 8 so you got time x) Alright #4 I'm gonna make you do over again and let me know what you get ;P

OpenStudy (anonymous):

lmao so in o.O I'm hungry. Btw I'm working on it right now.

OpenStudy (johnweldon1993):

I am too, that's why dinner is perfect Good food, awesome company...I'm golden XD

OpenStudy (anonymous):

Oh you're too sweet lol xP So I got f(x) = (x+2)^2-5 but I'm not confident about my answer :/

OpenStudy (johnweldon1993):

What if I told you the fate of dinner was relying on the confidence of that answer? xP Because you are absolutely Correct!!!!! :DDDD :D :D <3 XD

OpenStudy (anonymous):

OMFG I cannot believe it!! I literally read the first thing and was like "oh no dinner for me" xP Yay omg thank you so much!

OpenStudy (johnweldon1993):

Haha nope you're perfect!!! :D lol have more confidence in your answers though, you're very smart :)

OpenStudy (anonymous):

That's truly so sweet of you thank you! :]

OpenStudy (johnweldon1993):

Of course Alright, I must head to work now I'll pick ya up at 8 x) <3 Lol nice to meet you though, have a good day :)

OpenStudy (anonymous):

Haha! Thank you so much and it was a pleasure meeting you too. Good luck with work! :] Bye, you have a good day too.